# Two-Stage Weighted Projection for Reliable Low-Complexity Cooperative and Non-Cooperative Localization

### v1

@@ Abstract
Index Terms:
Two-stage weighted projection, low complexity positioning, cooperative localization, NLOS bias, TDOA, TOA, CRLB, two-way ranging.
@@ I Introduction
Localization is an essential component for numerous applications spanning navigation, autonomous systems, the Internet of Things (IoT), and next-generation communication networks like 5G Advanced and 6G. Over the years, significant advancements have been made in positioning methodologies, leveraging techniques such as TOA and TDOA-based localization. Among these, sophisticated methods such as nonlinear least squares (NLS) [2] and weighted NLS (WNLS) [3], along with optimization techniques like semi-definite programming (SDP) [4, 5] and multidimensional scaling (MDS) [6, 7] and machine learning-based techniques [8, 9] are recognized for their increased accuracy. However, they tend to be computationally intensive. In contrast, parallel projection methods (PPMs) [10, 11] are noted for their computational efficiency by avoiding complex operations like matrix inversion but are inferior to methods like NLS [12]. Note that PPMs have traditionally been confined to TOA-based localization, while SDP, MDS, and NLS methods are applicable to both TOA and TDOA-based localization. TOA-based localization requires precise time synchronization between anchors and user equipment (UEs), as well as awareness of signal transmission time. Conversely, TDOA-based localization does not require precise synchronization between anchors and UEs, is more resilient to clock offsets, and is more robust to multipath propagation. As a result, TDOA-based localization is often preferred, making methods like NLS, WNLS, SDP, and MDS often desirable. However, these methods may not be computationally feasible, especially for scenarios such as edge location computing (e.g., augmented reality (AR)/virtual reality (VR), drone navigation, and emergency geolocation) [13, 14], UE-driven non-terrestrial networks (NTN) access [15, 16], and power- and bandwidth-limited devices [17]. Motivated by these challenges, this paper proposes a novel two-stage weighted projection method (TS-WPM) for cooperative localization that provides provable gains in positioning accuracy while maintaining computational efficiency. Our proposed approach is particularly effective in challenging scenarios such as low SNR or high GDOP environments.
@@ I-A Related Work
Localization algorithms have been extensively studied in the literature, with numerous efforts dedicated to developing low-complexity methods that enhance accuracy while maintaining computational efficiency. Specifically, the work in [18] introduced a low-complexity optimization technique for localization using successive greedy grid search and successive cancellation-list grid search to address a measurement-to-target association problem, while [19] provides a framework for a low-complexity least-squares algorithm that demonstrates improved accuracy; however, their performance remains comparable to NLS and falls short of more advanced approaches such as WNLS and optimization-based methods. Furthermore, in  [20, 21], the authors introduced approaches for weighted PPM to enable enhanced cooperative localization, where the weights are determined using the gradient descent technique and the covariance of TOA measurements respectively. While these methods demonstrated promising results for TOA-based localization, they cannot be directly applied to TDOA-based scenarios. Further, they cannot be used if the TOA values are pseudo-range measurements, meaning they inherently include clock offsets and bias terms. In scenarios where there is no prior information on transmission time and bias offsets, such as clock offsets and non-line-of-sight (NLOS) bias, TOA values cannot be directly treated as noisy range measurements. Addressing these challenges, the authors in [12] proposed a reformulated version of IPPM to extend its applicability to TDOA-based localization. However, it is not straightforward to introduce suitable weighting mechanisms in PPMs to prevent overemphasizing or under-emphasizing specific TDOA measurements, particularly when certain anchor-to-UE links experience low SNR [1]. To overcome these limitations and achieve a balance between improved accuracy and computational feasibility, the current paper introduces a novel TS-WPM for robust cooperative localization. The proposed approach significantly enhances positioning accuracy, outperforming state-of-the-art approaches, including IPPM, NLS, and WNLS, while maintaining efficiency suitable for practical deployment.
@@ II-A Novel TS-WPM: TDOA Localization
To achieve a balance between improved accuracy and maintaining computational efficiency, we introduce a novel TS-WPM designed for TDOA-based localization. This new approach is particularly crucial for addressing the challenges associated with conditions such as high GDOP or low SNR regime. In this approach, we introduce additional variables to the estimation along with the target UE location. Specifically, we estimate the range or TOA between the reference anchor and the target UE, iteratively updating this estimate using the latest UE location. This updated range estimate is then utilized to refine the UE location estimate further.
@@
$\small\displaystyle\tilde{r}_{b}=\hat{r}_{b}-\hat{r}_{1},\quad b\in\{2,3,\cdots,B\}.$ (1)
The parameter vector to be estimated, denoted by $\small\boldsymbol{\theta}$, comprises the target UE coordinates. Denoting an estimate of the location by $\small\hat{\boldsymbol{\theta}}$, as part of TDOA-localization, we aim to minimize the TDOA residual error calculated as
where $\small\tilde{r}_{b,e}(\hat{\boldsymbol{\theta}})$ represents an estimate of TDOA associated with the anchor $\small b$ calculated as
$\small\displaystyle\tilde{r}_{b,e}(\hat{\boldsymbol{\theta}})=\|\hat{\boldsymbol{\theta}}-\mathbf{a}_{b}\|-\|\hat{\boldsymbol{\theta}}-\mathbf{a}_{1}\|,$ (3)
where $\small\mathbf{a}_{b}$ denotes the coordinates of the anchor $\small b$.
In our proposed approach, we introduce an additional parameter $\small r_{1}$ in the estimation process alongside $\small\boldsymbol{\theta}$, where $\small r_{1}$ enables an accurate estimation of the range between the reference anchor and the UE, thereby reducing TDOA residual errors and improving the localization accuracy. To facilitate this, we define a new variable
@@
where $\small\hat{r}_{1}^{e}$ is an estimate of $\small r_{1}$, and unlike $\small\hat{r}_{1}$, which is a pseudo-range, $\small\hat{r}_{1}^{e}$ is intended to be an accurate estimate of the true range. Importantly, we are not modifying the actual TDOA measurements but rather introducing a new variable that serves as the basis for the proposed approach and the corresponding analytical study. The proposed method still uses conventional TDOA measurements to update UE location, residual error, and range estimates, but the analyses and refinement process are guided by $\small\tilde{r}^{\prime}_{b}$. Given $\small\hat{r}_{1}^{e}$, $\small\tilde{r}^{\prime}_{b}$ become independent across anchors resembling TOA-like measurements. This ensures that, given $\small\hat{r}_{1}^{e}$, $\small{\rm var}(\tilde{r}^{\prime}_{b})$ is essentially equal to $\small{\rm var}(\hat{r}_{b})$. To incorporate this refinement into the estimation process, we redefine the residual error to be minimized as
@@
$\small\mathbf{R}^{0}\leftarrow\frac{1}{B}\sum_{b=1}^{B}\left(\tilde{r}_{b}-\big{(}\|\hat{\boldsymbol{\theta}}^{0}-\mathbf{a}_{b}\|-\hat{r}_{1}^{e,0}\big{)}\right)^{2}$
@@ 20       $\small k\leftarrow k+1$
The proposed novel TS-WPM is detailed in Algorithm 1, where weights are applied during the updates of both the UE location estimates and the range estimates. As illustrated, the algorithm starts by initializing the UE location estimate, denoted by $\small\hat{\boldsymbol{\theta}}^{0}$, which is usually assigned the coordinates of the nearest anchor, and the initial range estimate associated with the reference anchor, denoted by $\small\hat{r}_{1}^{e,0}$, which is usually set to 0. At the $\small k$-th iteration, the UE location estimate, $\small\hat{\boldsymbol{\theta}}^{k}$, is iteratively updated in step 1 using the previous location estimate, $\small\hat{\boldsymbol{\theta}}^{k-1}$, along with the refined range estimate, $\small\hat{r}_{1}^{e,k-1}$, obtained from iteration $\small k-1$. Concurrently, the TDOA residual error is updated in step 1 using the position estimate at the $\small k$-th iteration and the range estimate from the previous iteration. This range estimate is iteratively updated and accurately estimated in step 1, which, in turn, refines the UE location estimate in step 1 in the following iterations. This two-stage iterative process, by effectively mitigating the impact of bias offsets, including NLOS bias, leads to a continuous enhancement of positioning accuracy. This approach is particularly advantageous in scenarios with significant bias offsets, e.g., large NLOS bias and low SNR, where traditional methods struggle to maintain accuracy. The two-stage iterative refinement in TS-WPM has a conceptual resemblance to Turbo decoding principles [25], where extrinsic information is exchanged between decoders to improve error correction. Similarly, in TS-WPM, refining range estimates and UE location updates iteratively improves positioning accuracy. However, unlike Turbo decoding, where exchanged information is independent in terms of redundant parity, the refined information in TS-WPM is correlated, as both updates are derived from the same set of TDOA measurements.
@@ III-A Key Preliminaries
$\small\displaystyle\mathbf{F}_{2,[x,y]}=\mathbf{H}_{t}^{T}\mathbf{C}_{2}^{-1}\mathbf{H}_{t}-\mathbf{H}_{t}^{T}\mathbf{C}_{2}^{-1}\mathbf{v}\left(\mathbf{v}^{T}\mathbf{C}_{2}^{-1}\mathbf{v}\right)^{-1}\mathbf{v}^{T}\mathbf{C}_{2}^{-1}\mathbf{H}_{t},$ (17)
where $\small\mathbf{H}_{t}^{T}\mathbf{C}_{2}^{-1}\mathbf{v}\left(\mathbf{v}^{T}\mathbf{C}_{2}^{-1}\mathbf{v}\right)^{-1}\mathbf{v}^{T}\mathbf{C}_{2}^{-1}\mathbf{H}_{t}$ is the loss of information introduced due to estimating $\small r_{1}$. Not surprisingly, both formulations in (1) and (4) contain the same amount of information, i.e., $\small\mathbf{F}_{1}=\mathbf{F}_{2,[x,y]}$, resulting in identical CRLB.
@@ IV-A CRLB Analysis: CIR and AWGN
In this section, we derive the CRLB for TOA error in the presence of multipath propagation. As a special case of this, we also derive the CRLB for TOA error under AWGN. Let $\small\mathcal{N}_{b}=\{1,2,\ldots,B\}$ represent the set of anchors assigned to a target UE for PRS transmission. The time domain orthogonal frequency division multiplexing (OFDM) received signal between anchor $\small b$ and the target UE can be represented in matrix form as
@@ V Results and Discussion
In this section, we evaluate the positioning performance of the proposed TS-WPM and benchmark it against CRLB. We begin by outlining the evaluation framework. Our evaluation considers the anchor-UE deployment scenario illustrated in Fig. 2, where UEs are uniformly distributed within a predefined rectangular area, while anchors are strategically positioned around this area to optimize coverage and geometric diversity. In this configuration, each UE attempts to measure TOA from all visible anchors and subsequently computes TDOA measurements. TDOA measurements are computed using the anchor with the strongest SNR as the reference anchor. For this study, we assume that all UEs are located indoors, leading to predominantly NLOS conditions between the anchors and UEs. This aligns with 3GPP evaluation scenarios [24, 23] and avoids biases that could arise from scenarios favoring higher accuracy. The dual-stripe path loss model [23] is employed to account for distance-dependent path loss and penetration losses. To capture realistic channel characteristics, shadow fading effects are also considered [23]. To incorporate small-scale fading, we utilize clustered delay line (CDL) channel models, specifically the CDL-A channel model, which is well-suited for NLOS scenarios [24]. The system operates at a carrier frequency of 3.5 GHz with a 5 MHz bandwidth. We assume a maximum transmit (Tx) power of 23 dBm. This study demonstrates the effectiveness of our proposed TS-WPM by comparing it with state-of-the-art approaches including IPPM, NLS, and WNLS, with all results benchmarked against the CRLB. Additionally, we present a computational complexity analysis to compare the efficiency of TS-WPM with these state-of-the-art approaches. We specifically selected IPPM, NLS, and WNLS as comparison baselines for the following reasons: 1) The proposed TS-WPM is an iterative approach that updates location estimates in a manner similar to IPPM, NLS, and WNLS methods, making these comparisons particularly relevant. 2) Optimization-based techniques such as SDP [4, 5] and MDS [6, 7] are known to outperform weighted least squares (WLS) but offer performance comparable to NLS and are inferior to WNLS, making both NLS methods appropriate benchmarks.

### v3

@@ I Introduction
Localization is an essential component for numerous applications spanning navigation, autonomous systems, the Internet of Things (IoT), and next-generation communication networks like 5G Advanced and 6G. Over the years, significant advancements have been made in positioning methodologies, leveraging techniques such as TOA and TDOA-based localization. Among these, sophisticated methods such as nonlinear least squares (NLS) [2] and weighted NLS (WNLS) [3], along with optimization techniques like semi-definite programming (SDP) [4, 5, 6] and multidimensional scaling (MDS) [7, 8] and machine learning-based techniques [9, 10] are recognized for their increased accuracy. However, they tend to be computationally intensive. In contrast, parallel projection methods (PPMs) [11, 12] are noted for their computational efficiency by avoiding complex operations like matrix inversion but are inferior to methods like NLS [13]. Note that PPMs have traditionally been confined to TOA-based localization, while SDP, MDS, and NLS methods are applicable to both TOA and TDOA-based localization. TOA-based localization requires precise time synchronization between anchors and user equipment (UEs), as well as awareness of signal transmission time. Conversely, TDOA-based localization does not require precise synchronization between anchors and UEs, is more resilient to clock offsets, and is more robust to multipath propagation. As a result, TDOA-based localization is often preferred, making methods like NLS, WNLS, SDP, and MDS often desirable. However, these methods may not be computationally feasible, especially for scenarios such as edge location computing (e.g., augmented reality (AR)/virtual reality (VR), drone navigation, and emergency geolocation) [14, 15], UE-driven non-terrestrial networks (NTN) access [16, 17], and power- and bandwidth-limited devices [18]. Motivated by these challenges, this paper proposes a novel two-stage weighted projection method (TS-WPM) for cooperative localization that provides provable gains in positioning accuracy while maintaining computational efficiency. Our proposed approach is particularly effective in challenging scenarios, such as low SNR or high GDOP environments (see Appendix -A for the detailed definition of SNR and GDOP and the corresponding metrics used for evaluation purposes in this article).
@@ I-A Related Work
TABLE I: Comparison of Related Work with the Proposed TS-WPM Approach
Reference Key Contributions Scope Limitations / Gaps
Zhuo et al. [19] Low-complexity TOA localization using greedy and list-based grid search methods TOA-only Heuristic approach with significant gap from CRLB; no support for cooperative localization
Wu et al. [20] LS-based closed-form localization with improved stability under noise TOA-only No support for TDOA; does not account for SNR or multipath; no cooperative localization; lacks CRLB-based comparison
Liu et al. [21] Weighted PPM for TOA localization using SNR-based gradient descent optimization TOA-only Not applicable to TDOA; requires synchronization between anchors and UEs; limited positioning accuracy; no cooperative localization
Liu et al. [22] Covariance-aware weighted PPM in cooperative settings TOA-only No support for TDOA; lacks SNR or geometry-aware weighting; assumes tight anchor-UE synchronization; moderate accuracy
Dureppagari et al. [13] Reformulated IPPM for TDOA with a 3GPP-compliant CRLB framework under multipath TOA & TDOA No weighting strategy; moderate accuracy; inferior to or best matches with NLS
This Work Two-stage refinement with high accuracy, low complexity, robustness under high GDOP and low SNR, near-optimal performance at high SNR, and 3GPP-compliant evaluation TOA & TDOA
@@
Localization algorithms have been extensively studied in the literature, with numerous efforts dedicated to developing low-complexity methods that enhance accuracy while maintaining computational efficiency. Specifically, the work in [19] introduced a low-complexity optimization technique for localization using successive greedy grid search and successive cancellation-list grid search to address a measurement-to-target association problem, while [20] provides a framework for a low-complexity least-squares algorithm that demonstrates improved accuracy; however, their performance remains comparable to NLS and falls short of more advanced approaches such as WNLS and optimization-based methods. Furthermore, in  [21, 22], the authors introduced approaches for weighted PPM to enable enhanced cooperative localization, where the weights are determined using the gradient descent technique and the covariance of TOA measurements, respectively. While these methods demonstrated promising results for TOA-based localization, they cannot be directly applied to TDOA-based scenarios. Further, they cannot be used if the TOA values are pseudorange measurements, meaning they inherently include clock offsets and bias terms. In scenarios where there is no prior information on transmission time and bias offsets, such as clock offsets and NLOS bias, TOA values cannot be directly treated as noisy range measurements. Addressing these challenges, in our earlier work [13], we proposed a reformulated version of IPPM to extend its applicability to TDOA-based localization. Additionally, we introduced a 3GPP-compliant analytical framework using the CRLB tool that accounts for 5G NR physical layer parameters and multipath fading effects, which we extend in this article for evaluating the proposed TS-WPM. However, it is not straightforward to introduce suitable weighting mechanisms in PPMs to prevent overemphasizing or under-emphasizing specific TDOA measurements, particularly when certain anchor-to-UE links experience low SNR [1]. To overcome these limitations and achieve a balance between improved accuracy and computational feasibility, the current paper introduces a novel TS-WPM for robust cooperative localization. The proposed approach significantly enhances positioning accuracy, outperforming state-of-the-art approaches, including IPPM, NLS, and WNLS, while maintaining efficiency suitable for practical deployment. To clearly highlight the current research landscape and position our contribution, we present a comparative summary in Table I, which outlines the scope, strengths, and limitations of representative prior works in relation to TS-WPM.
@@ II-A Novel TS-WPM: TDOA Localization
To achieve a balance between improved accuracy and maintaining computational efficiency, we introduce a novel TS-WPM designed for TDOA-based localization. This new approach is particularly crucial for addressing the challenges associated with conditions such as high GDOP or low SNR regimes. In this approach, we introduce additional variables to the estimation along with the target UE location. Specifically, we estimate the range or TOA between the reference anchor and the target UE, iteratively updating this estimate using the latest UE location. This updated range estimate is then utilized to refine the UE location estimate further.
Consider a network of $\small B$ anchors assisting a target UE to obtain pseudorange or TOA measurements. The pseudorange measurements between an anchor $\small b$ and a target UE, denoted by $\small\hat{r}_{b}$, are generally defined as
$\small\displaystyle\hat{r}_{b}=\|\boldsymbol{\theta}-\mathbf{a}_{b}\|+\Delta_{b}+v_{b},\quad b\in\{1,\cdots,B\}.$ (1)
where $\small\mathbf{a}_{b}$ denotes known location coordinates of the anchor $\small b$, $\small\boldsymbol{\theta}$ is the true UE location, $\small\|\boldsymbol{\theta}-\mathbf{a}_{b}\|$ is the range between the anchor $\small b$ and the UE, $\small\Delta_{b}$ is the bias induced by multipath, and $\small n_{b}$ is additive white Gaussian noise (AWGN). Assuming the pseudorange measurements $\small\hat{r}_{b}$ are statistically independent across anchors, the covariance matrix of the TOA measurement vector $\small\hat{\mathbf{r}}=[\hat{r}_{1},\cdots,\hat{r}_{B}]^{T}$ is diagonal and takes the form $\small\hat{\boldsymbol{\Sigma}}=\mathrm{diag}(\sigma_{1}^{2},\cdots,\sigma_{B}^{2})$, where $\small\sigma_{b}^{2}$ is the variance of $\small v_{b}$.
Now, without loss of generality, we assume anchor $\small 1$ as the reference anchor. The TDOA measurements are then defined as
$\small\displaystyle\tilde{r}_{b}=\hat{r}_{b}-\hat{r}_{1},\quad b\in\{2,\dots,B\},$ (2)
Typically, the anchor with the strongest SNR is selected as the reference anchor [3]. This choice minimizes the variance of the common reference term $\small\hat{r}_{1}$, thereby reducing the overall correlation among TDOA measurements and improving localization robustness. Unlike pseudorange measurements, TDOA measurements share the common reference term $\small\hat{r}_{1}$, making them mutually dependent. As a result, the covariance matrix of the TDOA measurement vector $\small\tilde{\mathbf{r}}=[\tilde{r}_{2},\ldots,\tilde{r}_{B}]^{T}$ becomes non-diagonal. It takes the form $\small\tilde{\boldsymbol{\Sigma}}=\mathrm{diag}([\sigma_{2}^{2},\dots,\sigma_{B}^{2}])+\sigma_{1}^{2}\mathbf{u}\mathbf{u}^{T}$, where $\small\mathbf{u}$ is a $\small(B-1)\times 1$ vector of ones, capturing the correlation induced by the common reference measurement $\small\hat{r}_{1}$. Denoting an estimate of the UE location by $\small\hat{\boldsymbol{\theta}}$, as part of TDOA-localization, we aim to minimize the TDOA residual error calculated as
@@
$\small\displaystyle\tilde{r}_{b,e}(\hat{\boldsymbol{\theta}})=\|\hat{\boldsymbol{\theta}}-\mathbf{a}_{b}\|-\|\hat{\boldsymbol{\theta}}-\mathbf{a}_{1}\|.$ (4)
In our proposed approach, we introduce an additional unknown parameter $\small r_{1}$ in the estimation process alongside $\small\boldsymbol{\theta}$, where $\small r_{1}$ enables an accurate estimation of the range between the reference anchor and the UE, thereby reducing TDOA residual errors and improving the localization accuracy. To facilitate this, we define a new variable
@@
where $\small\hat{r}_{1}^{\,e}$ denotes an estimate of true range from the reference anchor to the UE. Unlike the pseudorange $\small\hat{r}_{1}$, which includes unknown clock biases and is inherently noisy, $\small\hat{r}_{1}^{\,e}$ is iteratively refined to approach the true (bias-free) range. Importantly, we are not modifying the actual TDOA measurements but rather introducing a new variable that serves as the basis for the proposed approach and the corresponding analytical study. The proposed method still uses conventional TDOA measurements to update UE location, residual error, and range estimates, but the analyses and refinement process are guided by $\small\tilde{r}^{\prime}_{b}$. Given $\small\hat{r}_{1}^{e}$, $\small\tilde{r}^{\prime}_{b}$ become independent across anchors resembling TOA-like measurements. This ensures that, given $\small\hat{r}_{1}^{e}$, $\small{\rm var}(\tilde{r}^{\prime}_{b})$ is essentially equal to $\small{\rm var}(\hat{r}_{b})$. To incorporate this refinement into the estimation process, we redefine the residual error to be minimized as
@@
[…]
@@ 20  $\small k\leftarrow k+1$
The proposed novel TS-WPM is detailed in Algorithm 1, where weights are applied during the updates of both the UE location estimates and the range estimates. As illustrated, the algorithm starts by initializing the UE location estimate, denoted by $\small\hat{\boldsymbol{\theta}}^{0}$, which is usually assigned the coordinates of the nearest anchor, and the initial range estimate associated with the reference anchor, denoted by $\small\hat{r}_{1}^{e,0}$ is set to $\small\|\hat{\boldsymbol{\theta}}^{0}-\mathbf{a}_{1}\|$ or the actual range estimate $\small\hat{r}_{1}$ if operating in high SNR regime. At the $\small k$-th iteration, the UE location estimate, $\small\hat{\boldsymbol{\theta}}^{k}$, is iteratively updated in step 1 using the previous location estimate, $\small\hat{\boldsymbol{\theta}}^{k-1}$, along with the refined range estimate, $\small\hat{r}_{1}^{e,k-1}$, obtained from iteration $\small k-1$. Concurrently, the TDOA residual error is updated in step 1 using the position estimate at the $\small k$-th iteration and the range estimate from the previous iteration. This range estimate is iteratively updated and accurately estimated in step 1, which, in turn, refines the UE location estimate in step 1 in the following iterations. This two-stage iterative process, by effectively mitigating the impact of bias offsets, including NLOS bias, leads to a continuous enhancement of positioning accuracy. This approach is particularly advantageous in scenarios with significant bias offsets, e.g., large NLOS bias and low SNR, where traditional methods struggle to maintain accuracy. The two-stage iterative refinement in TS-WPM has a conceptual resemblance to Turbo decoding principles [26], where extrinsic information is exchanged between decoders to improve error correction. Similarly, in TS-WPM, refining range estimates and UE location updates iteratively improves positioning accuracy. However, unlike Turbo decoding, where exchanged information is independent in terms of redundant parity, the refined information in TS-WPM is correlated, as both updates are derived from the same set of TDOA measurements.
@@ III-A Key Preliminaries
$\small\displaystyle\mathbf{F}_{2,[x,y]}=\mathbf{H}_{t}^{T}\mathbf{C}_{2}^{-1}\mathbf{H}_{t}-\mathbf{H}_{t}^{T}\mathbf{C}_{2}^{-1}\mathbf{v}\left(\mathbf{v}^{T}\mathbf{C}_{2}^{-1}\mathbf{v}\right)^{-1}\mathbf{v}^{T}\mathbf{C}_{2}^{-1}\mathbf{H}_{t}.$ (18)
The second term in (18) quantifies the loss of Fisher information resulting from the estimation of the additional parameter $\small r_{1}$. When additional (nuisance) parameters are jointly estimated along with the parameters of interest (in this case, UE position), the effective information available for position estimation is reduced. The term (18) explicitly captures the amount of information “spent” on estimating $\small r_{1}$, thereby quantifying the degradation of localization performance compared to a scenario where $\small r_{1}$ is known a priori. Not surprisingly, both formulations in (2) and (5) contain the same amount of information, i.e., $\small\mathbf{F}_{1}=\mathbf{F}_{2,[x,y]}$, resulting in identical CRLB.
@@ IV-A CRLB Analysis: CIR and AWGN
In this section, we derive the CRLB for TOA error in the presence of multipath propagation. As a special case of this, we also derive the CRLB for TOA error under AWGN. Let $\small\mathcal{N}_{b}=\{1,2,\ldots,B\}$ represent the set of anchors assigned to a target UE for PRS transmission. We adopt an orthogonal frequency division multiplexing (OFDM)-based received signal model, as 5G NR employs OFDM as its primary waveform [23]. However, OFDM-specific impairments such as timing and frequency synchronization errors, phase noise, and high peak-to-average power ratio (PAPR) are not explicitly modeled, as they fall outside the scope of this work. This is because we assume that TOA measurements are already available, and the received signal model is used solely for deriving CRLB expressions, which in turn are used to formulate TOA error models for Monte Carlo simulations in evaluating the performance of the proposed TS-WPM algorithm. The time domain OFDM received signal between the anchor $\small b$ and the target UE can be represented in matrix form as
@@ V Results and Discussion
In this section, we evaluate the positioning performance of the proposed TS-WPM and benchmark it against CRLB. We begin by outlining the evaluation framework. Our evaluation considers the anchor-UE deployment scenario illustrated in Fig. 2, where UEs are uniformly distributed within a predefined rectangular area, while anchors are strategically positioned around this area to optimize coverage and geometric diversity. In this configuration, each UE attempts to measure TOA from all visible anchors and subsequently computes TDOA measurements. TDOA measurements are computed using the anchor with the strongest SNR as the reference anchor. For this study, we assume that all UEs are located indoors, leading to predominantly NLOS conditions between the anchors and UEs. This aligns with 3GPP evaluation scenarios [25, 24] and avoids biases that could arise from scenarios favoring higher accuracy. The dual-stripe path loss model [24] is employed to account for distance-dependent path loss and penetration losses. To capture realistic channel characteristics, shadow fading effects are also considered [24]. To incorporate small-scale fading, we utilize clustered delay line (CDL) channel models, specifically the CDL-A channel model, which is well-suited for NLOS scenarios [25]. The system operates at a carrier frequency of 3.5 GHz with a 5 MHz bandwidth. We assume a maximum transmit (Tx) power of 23 dBm. This study demonstrates the effectiveness of our proposed TS-WPM by comparing it with state-of-the-art approaches, including IPPM, NLS, and WNLS, with all results benchmarked against the CRLB. To provide a comprehensive performance benchmark, we also evaluate TS-WPM under an AWGN channel. This aligns with the closed-form CRLB expression derived in Corollary 2.1 and facilitates quantifying the performance degradation in multipath-rich environments by comparison with the generalized CRLB derived in Theorem 2. Additionally, we present a computational complexity analysis to compare the efficiency of TS-WPM with these state-of-the-art approaches.
The said baselines were chosen for the following reasons: 1) The proposed TS-WPM is an iterative approach that updates location estimates in a manner similar to IPPM, NLS, and WNLS methods, making these comparisons particularly relevant. 2) Optimization-based techniques such as SDP [4, 5, 6] and MDS [7, 8] are typically designed for TOA-only settings, offer moderate accuracy, are computationally complex, and do not offer optimal performance even under AWGN, unlike WNLS. 3) Machine learning-based methods [9, 10] require large datasets, extensive offline training, and lack analytical bounds. Moreover, environment-specific training/validation limits their generalizability, e.g., low SNR or high GDOP scenarios. 4) WNLS remains one of the most sophisticated and widely adopted estimators for TDOA-based localization, offering near-optimal performance under favorable conditions [36, 37]. Furthermore, several recent works [19, 21, 22] have explored low-complexity localization strategies including greedy search, weighted PPM, and covariance-aware refinement. While promising, these methods are limited to TOA-only settings and do not support TDOA, and evaluation under realistic conditions, such as multipath and radio propagation aspects. Thus, by considering WNLS as the strongest conventional benchmark and demonstrating the superiority and near-CRLB performance of TS-WPM across various conditions (e.g., high GDOP, low SNR, multipath), our evaluation implicitly establishes the robustness and practical relevance of TS-WPM, even when compared to the latest localization techniques. Note that the IPPM baseline included in this study is the reformulated version we proposed in our earlier work [13], where we extended traditional IPPM to support TDOA-based localization. As shown in [13], while this extended IPPM offers lower complexity, its performance is comparable to or matches that of NLS at best, a finding further reinforced through our results in this section.
Moreover, in all experiments, we set the convergence threshold to $\small\epsilon=10^{-7}$, and require the residual error to remain below this threshold for $\small l=10$ consecutive iterations before terminating the iterative loop and declaring convergence. While $\small\epsilon=10^{-7}$ enables high precision convergence, the choice of $\small l=10$ prevents premature termination due to transient noise, thereby ensuring convergence stability. Additionally, we put a hard limit on the total number of iterations at 100 to ensure computational feasibility and bounded complexity, especially under high-GDOP or low-SNR conditions. These values were determined through empirical validation across diverse deployment scenarios, providing a robust trade-off between accuracy and efficiency.